What is angular frequency.


1
Expert's answer
2021-02-01T08:03:55-0500

Angular frequency is the angular displacement per unit time. The SI unit of angular frequency is radian per second. The formula for angular frequency can be written as follows:


"\omega=\dfrac{2\pi}{T}=2\pi f."